1.)Calculate the volume, in liters, of 1.23 molH2S(g)at 61 ∘C and 1.30 atm.

2.) Calculate the pressure, in atmospheres, of 4.86 molCO(g)in a 3.70 Ltank at 27 ∘C.

Express your answer to three significant figures.
Answer #1

1)

Given:

P = 1.3 atm

n = 1.23 mol

T = 61.0 oC

= (61.0+273) K

= 334 K

use:

P * V = n*R*T

1.3 atm * V = 1.23 mol* 0.08206 atm.L/mol.K * 334 K

V = 25.9322 L

Answer: 25.9 L

2)

Given:

V = 3.7 L

n = 4.86 mol

T = 27.0 oC

= (27.0+273) K

= 300 K

use:

P * V = n*R*T

P * 3.7 L = 4.86 mol* 0.08206 atm.L/mol.K * 300 K

P = 32.3361 atm

Answer: 32.3 atm